OpenStudy (anonymous):

w=Cr^-2Solve the equation for r.

OpenStudy (anonymous):

Okay. This is a bit outside of the intuitive zone. Starting off, we will divide both sides of the equation by C. This gives us: w/C = r^-2 But r^-2 is the same as 1/r^2 So we have the following equation: w/C = 1/r^2 If we reciprocate both sides of the equation, we have: C/w = r^2 Take the square root of both sides, and you have: sqrt(C/w) = r Hope this helps! Visit my website at http://www.tutorsean.net for free math and computer science videos.